OpenStudy (help!!!!):

Mike weights x-3, his brother Tim weights 8 more pounds that mike. While his (Sarah)sister weights 12 pounds than tim. Find their total weight.

OpenStudy (help!!!!):

a.4x-5 b.5x+3 c.x+12 d.3x-5 e.2x+5

OpenStudy (anonymous):

3x+19 is the answer .. question might be typed wrong .. plzz look into it again

OpenStudy (help!!!!):

thats what i got, but the answer is d

OpenStudy (help!!!!):

i dont understand

OpenStudy (anonymous):

got it sarah is 12 pounds less than tim then Mike = x-3 tim = x-3+8 = x+5 sarah = x+5-12 = x-7 total = 3x -5